if a caterer charges a fixed cost for preparing a dinner plus an additional cost for each person served and you know that the cos will be for 100 students will be $750 and the cost for 150 students will be $1050,what would the caterer's fixed cost and the cost per student be?(using linear equations)
Answer by stanbon(75887)

if a caterer charges a fixed cost for preparing a dinner plus an additional cost for each person served and you know that the cost will be for 100 students will be $750 and the cost for 150 students will be $1050,what would the caterer's fixed cost and the cost per student be?(using linear equations)
------
You have two points relating # of students and cost
(100,750) and (150,1050)
---
slope = (1050-750)/(150-100) = 300/50 = 6
---
Form: c = mx + b
750 = 6*100 + b
b = 150
----
Cost(x) = 6x + 150
---
Cost per student = $6
==========================
Cheers,
Stan H.